At standard temperature and pressure (0 ∘C and 1.00 bar), 1.00 mol of an ideal gas occupies a volume of 22.7 L. What volume would the same amount of gas occupy at the same pressure and 30 ∘C ?
1
Expert's answer
2018-01-31T05:12:08-0500
t(0)=0 С T=273 K
p=1.00 bar
V=22.7 l
N=1 mol
t(1)=30 С T=303 K
Find V=?
According to the Boyle-Mariotte law
P(0) V(0)/T(0) = const
If p=const
V(0)/T(0)= V(1)/T(1)
V(1)=V(0)T(1)/T(0)=22.7 *303/273=25.19 l